OpenStudy (anonymous):

Estimate by first rounding each number to the nearest 1/2. -3 8/9 * 1 /12 A. –6 B. –3 C. 6 D. 8

OpenStudy (anonymous):

@GretaKnows

OpenStudy (anonymous):

okay so 8/9 is 1/9 form being 4 so the first umber is 4. then.... wait is the second number 1/12?

OpenStudy (anonymous):

*wait i mean -4

OpenStudy (anonymous):

Yea one half

OpenStudy (anonymous):

one half or 1/12?

OpenStudy (anonymous):

One half

OpenStudy (anonymous):

okay so then the second number is 1/2.... -4*1/2=??

OpenStudy (anonymous):

but none of the answers are right then... are you sure you wrote the right numbers?

OpenStudy (anonymous):

Opps the 2nd one is 1/12

OpenStudy (anonymous):

I just looked back at my pappers

OpenStudy (anonymous):

okay so its -3 8/9 * 1/12?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

there is still not a right answer -3 8/9 rounds down to -3 1/2 and 1/12 rounds up to 1/2. -3 1/2 * 1/2 is -3/4 and thats not an answer. are you sure you arent supposed to add and not multiply?

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

okay well then i cant help you -3 1/2 * 1/2 is not any of the answers sorry i couldn't help you with this one...

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Godlovesme @GodGirl360

OpenStudy (danjs):

-3 8/9 to the closest half is , -4 1/12 to the nearest half is, 0, -4*0 = 0, that is not an answer

OpenStudy (danjs):

if you say 1/12 is rounded to 1/2 instead -4 * 1/2 = -2 that still isnt an answer

OpenStudy (danjs):

is the * supposed to be a + maybe

OpenStudy (anonymous):

no it's a *

OpenStudy (anonymous):

Yeah thats what I thought at first bu that isnt an answer then i tried to the nearest half as in like numbers with halfs in them like 3 1/2 but that didnt work either.., either they dont have the right answers or somebody wrote the problem wrong.... i already asked she said its multiplication.

OpenStudy (anonymous):

It's 1 1/2 one and one half

OpenStudy (danjs):

-4 * 1.5 = -6

OpenStudy (danjs):

-4 * 1 and a half = -6

OpenStudy (danjs):

if that is what you mean

OpenStudy (anonymous):

YES and i got 3 more

OpenStudy (danjs):

dang, make sure it is typed right :-P

OpenStudy (anonymous):

haha that confused me...

OpenStudy (danjs):

right, you have to interpret what they really meant to ask sometimes. lol

OpenStudy (anonymous):

i was like, do I not know how to multiply?!?! I thought I learned to do that...